OpenStudy (anonymous):

what is the least common demonator? for both(x^2-5x-6) and (x^2-4x-5)

OpenStudy (anonymous):

x+1 because (x-6)(x+1) = x^2-5x-6 and (x-5)(x+1) = x^2-4x-5 i.e. x+1 would divide into both of these